# Additive subtraction games

### 11 Mar 2026 v1

@@ 4. The arithmetic structure of the $\delta$-collisions
Recall that in this regime we assume $a<\delta<2a,\gcd(a,\delta)=1,d:=\delta-a$, and we let
$w_{n}=n+a\Bigl\lfloor\frac{n}{a}\Bigr\rfloor+(a+\delta)\Bigl\lfloor\frac{n}{\delta}\Bigr\rfloor.\qquad n\geqslant 0,$
We write $\mathcal{W}_{0}=\{w_{n}:n\geqslant 0\}$ as before. For a subset $X\subset\mathbb{N}_{0}$, we write $|X|_{p}$ for the number of elements of $X$ whose heaps lie in a single heap-period; in the primitive quadratic case this is the same as counting indices in a single index-period of length $a\delta$.
We are interested in the intersections of $\mathcal{W}_{0}$ with its $\delta$-shift $\mathcal{W}_{0}-\delta=\{w_{n}-\delta:n\geqslant 0\}$. Equivalently, we count all pairs $(m,n)$ with $n>m$ such that $w_{n}-w_{m}=\delta$.
@@
(11) $C:=\{\,w_{n}:\exists m<n\text{ with }w_{n}-w_{m}=\delta\,\}=(W_{0}-\delta)\cap W_{0}.$
@@
Assume $a<\delta<2a$ and $\gcd(a,\delta)=1$. Then, in one heap-period of length $(3\delta+a)a$, the number of collisions is $|C|_{p}=ad$.
@@ 4.1. Gaps and residue classes
The arithmetic structure of the $\delta$-collisions relies on a regularity property of the gap sequence of the candidate $\mathcal{P}$-positions. Recall
$w_{n}=n+a\Big\lfloor\frac{n}{a}\Big\rfloor+(a+\delta)\Big\lfloor\frac{n}{\delta}\Big\rfloor,\qquad n\geqslant 0.$
Define the gap sequence
@@
The key observation is that the gap sequence has a very simple arithmetic structure. Since the definition of $w_{n}$ involves only the two floor functions $\lfloor n/a\rfloor$ and $\lfloor n/\delta\rfloor$, the gap $g_{n}=w_{n+1}-w_{n}$ can change only when $n+1$ crosses a multiple of $a$ or a multiple of $\delta$. Consequently the gap sequence takes only four possible values and is completely determined by the residue classes of $n+1$ modulo $a$ and $\delta$. We record this observation explicitly.
@@
In particular the gap sequence is periodic with index-period $a\delta$, and the occurrences of the special gaps $a+1$ and $(a+\delta)+1$ are uniquely determined by the residue classes of $n+1$ modulo $a$ and $\delta$, respectively.
@@
$g_{n}=w_{n+1}-w_{n}=1+a\!\left(\Big\lfloor\frac{n+1}{a}\Big\rfloor-\Big\lfloor\frac{n}{a}\Big\rfloor\right)+(a+\delta)\!\left(\Big\lfloor\frac{n+1}{\delta}\Big\rfloor-\Big\lfloor\frac{n}{\delta}\Big\rfloor\right).$
@@
Let $n+1=qa$ and reduce modulo $\delta$. Set $r\equiv qa\pmod{\delta}$. Since $\gcd(a,\delta)=1$, multiplication by $a$ is invertible modulo $\delta$.
@@
As $q$ runs through $1,\dots,\delta-1$, the residues
$r\equiv qa\pmod{\delta}$
run bijectively through $\{1,\dots,\delta-1\}$. In particular, in one index-period of length $a\delta$ there are exactly $\delta-1$ gaps equal to $a+1$.
@@
Thus we can label each $(a+1)$-gap in a period by a unique residue $r\in\{1,\dots,\delta-1\}$.
@@ 4.2. Collision windows and the parameters $L$ and $R$
A collision window is a stretch of consecutive gaps whose sum equals $\delta=a+d$. Since $a<\delta<2a$, by (12) any such collision window must contain exactly one $(a+1)$-gap and exactly d1d-1 gaps of size one. Gaps greater than $\delta$ (i.e. $a+\delta$ and $2a+\delta$) cannot appear inside a collision window; in fact, they terminate any consecutive run of $1$-gaps. Fix one particular $(a+1)$-gap. Let $L$ and $R$ denote the numbers of consecutive $1$-gaps immediately to its left and right, respectively; for an example, see Figure 1.
@@
Let the $(a+1)$-gap correspond to residue $r\in\{1,\dots,\delta-1\}$ as above. Then
$L=\min(a-1,r-1),\qquad R=\min(a-1,\delta-r-1).$
@@
Between two successive multiples of $a$ there are exactly a1a-1 indices, and hence at most a1a-1 gaps of size $1$ appears before the next large gap. Thus, $L$ and $R$ will have size a1a-1 unless a larger gap appears before the next multiple of $a$.
Measuring from the nearest multiple of $\delta$ to the left, the residue $r$ counts how many steps one is away from that multiple. Thus there can be at most r1r-1 consecutive indices before reaching that multiple from the right, and at most $\delta-r-1$ indices before reaching the next multiple of $\delta$ to the right. Combining these two types of obstructions yields the stated minima. ∎
@@
Figure 1. Consider $a=4$ and $\delta=7$. The gaps along the number line have sizes $1,1,1,5,1,12$ where $12>\delta$. The central gap of size $a+1=5$ is flanked by $L=3$ consecutive one-gaps on the left and $R=1$ on the right; the large dashed gap $\delta$ prevents extending collision windows further right. The blue and red braces show the two possible collision windows of length $\delta=7$.
Let us denote by $c(r)$ the number of collision windows belonging to the $(a+1)$-gap at residue $r$. It must contain, $0\leqslant x\leqslant d-1$ one-gaps from the left and $0\leqslant d-1-x\leqslant d-1$ one-gaps from the right, because $a+1+x+d-1-x=\delta$. By Lemma 10, the admissible such $x$ values are those satisfying both
@@
4.3. Three residue regions and the distribution of $c(r)$
We now compute the number $c(r)$ of collision windows belonging to the $r$th $(a+1)$-gap explicitly. The numbers are computed by using the membership of the residue $r$ of the three relevant regions, two of which are restricted to the left and right, respectively. Recall that $r\equiv qa\pmod{\delta}$, where $q$ runs through $1,\ldots,\delta-1$, with $d=\delta-a$ and that we are in the quadratic regime where $d<a$.
Left-restricted: $1\leqslant r\leqslant d$. Here $r-1\leqslant d-1<a-1$, so
$L=\min(a-1,r-1)=r-1,\qquad R=\min(a-1,\delta-r-1)=a-1.$
Thus the right side does not restrict. The constraints (13) reduce to $0\leqslant x\leqslant r-1$. Hence $x$ can take the $r$ values $0,1,\dots,r-1$, and
(14) $c(r)=r.$
Unrestricted: $d<r<a$. Now $r-1\geqslant d$ and $\delta-r-1\geqslant d-1$, so both $L$ and $R$ are at least d1d-1, and the only restriction on $x$ comes from the requirement $0\leqslant d-1-x$. Hence $x$ can range freely from 0 to d1d-1, and
(15) $c(r)=d.$
Right-restricted: $a\leqslant r\leqslant\delta-1$. We have
$L$ $=\min(a-1,r-1)$
@@ 4.3. Three residue regions and the distribution of $c(r)$
$R$ $=\min(a-1,\delta-r-1)$
$=\delta-r-1,$
since $\delta-r-1\leqslant d-1<a-1$. Thus the constraints (13) become
$0\leqslant d-1-x\leqslant\delta-r-1.$
@@
Since $\gcd(a,\delta)=1$, multiplication by $a$ permutes the nonzero residues modulo $\delta$. Hence as $q$ runs through $1,\ldots,\delta-1$, the residues $r\equiv qa\pmod{\delta}$ run through the set $\{1,2,\ldots,\delta-1\}$ exactly once. Thus it suffices to sum $c(r)$ for $r=1,\ldots,\delta-1$. From the three cases above we have
$c(r)=\begin{cases}r,&1\leqslant r\leqslant d,\\ d,&d<r<a,\\ \delta-r,&a\leqslant r\leqslant\delta-1.\end{cases}$
In the left region ($1\leqslant r\leqslant d$) the values are $1,2,\ldots,d$. In the middle region ($d<r<a$) there are ad1a-d-1 residues and each contributes $d$. In the right region ($a\leqslant r\leqslant\delta-1$), the values $\delta-r$ run through $d,d-1,\ldots,1$. Thus the multiset of values is
$\{c(r):1\leqslant r\leqslant\delta-1\}=\{1,2,\ldots,d,\underbrace{d,\ldots,d}_{a-d-1\ \text{times}},d,\ldots,2,1\}.$
@@
rrxxaa
Figure 2. Each column $r$ contains exactly $c(r)$ admissible lattice points. Thus the total number of points equals the number of lattice points $(r,x)$ satisfying $r-a\leqslant x\leqslant r-1$ inside the rectangle $1\leqslant r\leqslant\delta-1$, $0\leqslant x\leqslant d-1$.
@@ 4.4. Proof of the $\delta$-collision counting theorem
In one index-period of length $a\delta$, the $(a+1)$-gaps are in bijection with residues $r\in\{1,\dots,\delta-1\}$, and for each $r$, $c(r)$ is the number of collision windows containing that gap. Thus the total number of collision windows (and therefore collisions) in one index-period equals
$\sum_{r=1}^{\delta-1}c(r).$
@@ 5. Some numerical examples
This section illustrates the arithmetic mechanism of Section 4 with three examples. The first shows the degenerate situation $d=1$, where every $(a+1)$-gap yields exactly one collision window. The second gives an explicit computation in the first non-trivial case $d=2$. The third demonstrates how the parameters $L$ and $R$ determine $c(r)$.
@@
The gaps of size $a+1=4$ occur exactly $\delta-1=3$ times. Since $d=1$, a collision window consists solely of a single $(a+1)$-gap and no additional $1$-gaps. Thus each such gap contributes exactly one collision window, so $c(r)=1$ for all $r$ and
$\sum_{r}c(r)=3=ad.$
@@
Since $d=2$, each collision window must contain one $(a+1)$-gap and one neighbouring $1$-gap. Inspecting the sequence shows that some six-gaps are adjacent to two $1$-gaps while others are adjacent to only one. Writing $c(r)$ for the number of collision windows containing the $(a+1)$-gap at index $r$, we obtain
@@
Hence among the $\delta-1=6$ many $(a+1)$-gaps, exactly two have $c(r)=1$ and the remaining four have $c(r)=2$. Therefore
$\sum_{r}c(r)=1+1+2+2+2+2=10=ad.$
@@
Let $a=4$ and $\delta=7$. Then $d=3$, so each collision window must contain exactly one gap of size $a+1=5$ and exactly d1=2d-1=2 gaps of size $1$. This example illustrates how the parameters $L$ and $R$ determine the number $c(r)$. The index-period has length $a\delta=28$. The gap formula reads
@@
Consider the gap corresponding to $n+1=12$. Writing $12=qa$ with $q=3$ and reducing modulo $\delta=7$ gives $12\equiv 5\pmod{7}$, so $r=5$. We now compute
$L=\min(a-1,r-1)=\min(3,4)=3,$
@@
force $x=1$ or $x=2$. Hence this $(a+1)$-gap contributes $c=2$ collision windows. This agrees with the right-restricted case of the general formula (where $a\leqslant r\leqslant a+d-1$).
6. Game theory conclusion
@@
In one heap-period, $|\mathcal{W}_{3}|_{p}=a^{2}$.
@@
By Theorem 7, we have that $|C|_{p}=ad$, where $d=\delta-a$ and consequently
$|\mathcal{W}_{3}|_{p}$ $=|W_{0}-\delta|_{p}-|C|_{p}$
@@
$\mathbb{N}_{0}=\mathcal{W}_{0}\sqcup\mathcal{W}_{1}\sqcup\mathcal{W}_{2}\sqcup\mathcal{W}_{3},$
@@
We have observed $|\mathcal{W}_{0}|_{p}=|\mathcal{W}_{1}|_{p}=|\mathcal{W}_{2}|_{p}=a\delta$, and, via Corollary 15, $|\mathcal{W}_{3}|_{p}=a^{2}$. In total, the disjoint union
@@
$|\mathcal{W}_{0}|_{p}+|\mathcal{W}_{1}|_{p}+|\mathcal{W}_{2}|_{p}+|\mathcal{W}_{3}|_{p}=3a\delta+a^{2}$
positions, exactly matching the heap-period length in the primitive quadratic regime. Recall, the index-period is $a\delta$, and since $a\delta$ is simultaneously a multiple of $a$ and $\delta$, the direct computation gives the corresponding period length
$w_{n+a\delta}-w_{n}=(3\delta+a)a.$

### 29 Mar 2026 v2

@@ 4. The arithmetic structure of the $\delta$-collisions
Recall that in this regime we assume $a<\delta<2a,\gcd(a,\delta)=1,d:=\delta-a$, and, for $n\geqslant 0$, we let
$w_{n}=n+a\Bigl\lfloor\frac{n}{a}\Bigr\rfloor+(a+\delta)\Bigl\lfloor\frac{n}{\delta}\Bigr\rfloor.$
We write $\mathcal{W}_{0}=\{w_{n}:n\geqslant 0\}$ as before. For a subset $X\subset\mathbb{N}_{0}$, we write $|X|_{h}$ for the number of elements of $X$ whose heaps lie in a single heap-period; in the primitive quadratic case this is the same as counting indices in a single index-period of length $a\delta$.
We are interested in the intersections of $\mathcal{W}_{0}$ with its $\delta$-shift $\mathcal{W}_{0}-\delta=\{w_{n}-\delta:n\geqslant a\}$. Equivalently, we count all pairs $(m,n)$ with $n>m$ such that $w_{n}-w_{m}=\delta$.
@@
(16) $C:=\{\,w_{n}:\exists m<n\text{ with }w_{n}-w_{m}=\delta\,\}=(\mathcal{W}_{0}-\delta)\cap\mathcal{W}_{0}.$
@@
Assume $a<\delta<2a$ and $\gcd(a,\delta)=1$. Then, in one heap-period of length $h=(3\delta+a)a$, the number of collisions is $|C|_{h}=ad$.
@@ 4.1. Gaps and residue classes
The arithmetic structure of the $\delta$-collisions relies on a regularity property of the gap sequence of the candidate $\mathcal{P}$-positions. Recall, for $n\geqslant 0$,
Define the gap sequence, for $n\geqslant 0$,
@@
The key observation is that the gap sequence has a very simple arithmetic structure. Since the definition of $w_{n}$ involves only the two floor functions $\lfloor n/a\rfloor$ and $\lfloor n/\delta\rfloor$, the gap $g_{n}$ can change only when $n+1$ crosses a multiple of $a$ or a multiple of $\delta$. Consequently the gap sequence takes only four possible values and is completely determined by the residue classes of $n+1$ modulo $a$ and $\delta$. We record this observation explicitly.
@@
In particular the gap sequence is periodic with index-period $a\delta$, and the occurrences of the special gaps $a+1$ and $b+1$ are uniquely determined by the residue classes of $n+1$ modulo $a$ and $\delta$, respectively.
@@
$g_{n}=w_{n+1}-w_{n}=1+a\!\left(\Big\lfloor\frac{n+1}{a}\Big\rfloor-\Big\lfloor\frac{n}{a}\Big\rfloor\right)+b\!\left(\Big\lfloor\frac{n+1}{\delta}\Big\rfloor-\Big\lfloor\frac{n}{\delta}\Big\rfloor\right).$
@@
Let $n+1=qa$ and reduce modulo $\delta$. Set $\rho\equiv qa\pmod{\delta}$. Since $\gcd(a,\delta)=1$, multiplication by $a$ is invertible modulo $\delta$.
@@
As $q$ runs through $1,\dots,\delta-1$, the residues $\rho\equiv qa\pmod{\delta}$ run bijectively through $\{1,\dots,\delta-1\}$. In particular, in one index-period of length $a\delta$ there are exactly $\delta-1$ gaps equal to $a+1$.
@@
Thus we can label each $(a+1)$-gap in a period by a unique residue $\rho\in\{1,\dots,\delta-1\}$.
@@ 4.2. Collision windows and the parameters $L$ and $R$
A collision window is a stretch of consecutive gaps whose sum equals $\delta=a+d$. Since $a<\delta<2a$, by (17) any such collision window must contain exactly one $(a+1)$-gap and exactly d1d-1 gaps of size one. Gaps greater than $\delta$ (i.e. $b+1$ and $a+b+1$) cannot appear inside a collision window; in fact, they terminate any consecutive run of $1$-gaps. Fix one particular $(a+1)$-gap. Let $L$ and $R$ denote the numbers of consecutive $1$-gaps immediately to its left and right, respectively; for an example, see Figure 3.
@@
Let the $(a+1)$-gap correspond to residue $\rho\in\{1,\dots,\delta-1\}$ as above. Then
$L=\min(a-1,\rho-1),\qquad R=\min(a-1,\delta-\rho-1).$
@@
Between two successive multiples of $a$ there are exactly a1a-1 indices, and hence at most a1a-1 gaps of size $1$ appear before the next large gap. Thus, $L$ and $R$ will have size a1a-1 unless a larger gap appears before the next multiple of $a$.
Measuring from the nearest multiple of $\delta$ to the left, the residue $\rho$ counts how many steps one is away from that multiple. Thus there can be at most $\rho-1$ consecutive indices before reaching that multiple from the right, and at most $\delta-\rho-1$ indices before reaching the next multiple of $\delta$ to the right. Combining these two types of obstructions yields the stated minima. ∎
@@
Figure 3. Consider $a=4$ and $\delta=7$. The gaps along the number line have sizes $1,1,1,5,1,12$ where $12>\delta$. The central gap of size $a+1=5$ is flanked by $L=3$ consecutive one-gaps on the left and $R=1$ on the right; the large dashed gap $\delta$” prevents extending collision windows further right. The blue and red braces show the two possible collision windows of length $\delta=7$.
Let us denote by $c(\rho)$ the number of collision windows belonging to the $(a+1)$-gap at residue $\rho$. It must contain, $0\leqslant x\leqslant d-1$ one-gaps from the left and $0\leqslant d-1-x\leqslant d-1$ one-gaps from the right, because $a+1+x+d-1-x=\delta$. By Lemma 12, the admissible such $x$ values are those satisfying both
@@
4.3. Three residue regions and the distribution of $c(\rho)$
[…]
Left-restricted: $1\leqslant\rho\leqslant d$. Here $\rho-1\leqslant d-1<a-1$, so
$L=\min(a-1,\rho-1)=\rho-1,\qquad R=\min(a-1,\delta-\rho-1)=a-1.$
Thus the right side does not restrict. The constraints (18) reduce to $0\leqslant x\leqslant\rho-1$. Hence $x$ can take the $\rho$ values $0,1,\dots,\rho-1$, and
(19) $c(\rho)=\rho.$
Unrestricted: $d<\rho<a$. Now $\rho-1\geqslant d$ and $\delta-\rho-1\geqslant d-1$, so both $L$ and $R$ are at least d1d-1, and the only restriction on $x$ comes from the requirement $0\leqslant d-1-x$. Hence $x$ can range freely from 0 to d1d-1, and
(20) $c(\rho)=d.$
Right-restricted: $a\leqslant\rho\leqslant\delta-1$. We have
$L$ $=\min(a-1,\rho-1)$
@@ 4.3. Three residue regions and the distribution of $c(\rho)$
$R$ $=\min(a-1,\delta-\rho-1)$
$=\delta-\rho-1,$
since $\delta-\rho-1\leqslant d-1<a-1$. Thus the constraints (18) become
$0\leqslant d-1-x\leqslant\delta-\rho-1.$
@@
Since $\gcd(a,\delta)=1$, multiplication by $a$ permutes the nonzero residues modulo $\delta$. Hence as $q$ runs through $1,\ldots,\delta-1$, the residues $\rho\equiv qa\pmod{\delta}$ run through the set $\{1,2,\ldots,\delta-1\}$ exactly once. Thus it suffices to sum $c(\rho)$ for $\rho=1,\ldots,\delta-1$. From the three cases above we have
$c(\rho)=\begin{cases}\rho,&1\leqslant\rho\leqslant d,\\ d,&d<\rho<a,\\ \delta-\rho,&a\leqslant\rho\leqslant\delta-1.\end{cases}$
In the left region ($1\leqslant\rho\leqslant d$) the values are $1,2,\ldots,d$. In the middle region ($d<\rho<a$) there are ad1a-d-1 residues and each contributes $d$. In the right region ($a\leqslant\rho\leqslant\delta-1$), the values $\delta-\rho$ run through $d,d-1,\ldots,1$. Thus the multiset of values is
$\{c(\rho):1\leqslant\rho\leqslant\delta-1\}=\{1,2,\ldots,d,\underbrace{d,\ldots,d}_{a-d-1\ \text{times}},d,\ldots,2,1\}.$
@@
[…]
Figure 4. The distribution of $c(\rho)$ across the three residue regions for $a<\delta<2a$ and $d=\delta-a$. Each column $\rho\in\{1,\dots,\delta-1\}$ contains exactly $c(\rho)$ admissible lattice points, where each point corresponds to a unique collision window containing the $(a+1)$-gap at residue $\rho$. The left triangle ($1\leqslant\rho\leqslant d$) represents the left-restricted case where $c(\rho)=\rho$; the central rectangle ($d<\rho<a$) represents the unrestricted case where $c(\rho)=d$; and the right triangle ($a\leqslant\rho\leqslant\delta-1$) represents the right-restricted case where $c(\rho)=\delta-\rho$. The total area (lattice point count) is $\sum_{\rho=1}^{\delta-1}c(\rho)=ad$.
@@ 4.4. Proof of the $\delta$-collision counting theorem
In one index-period of length $a\delta$, the $(a+1)$-gaps are in bijection with residues $\rho\in\{1,\dots,\delta-1\}$, and for each residue $\rho$, $c(\rho)$ is the number of collision windows containing that gap. Thus the total number of collision windows (and therefore collisions) in one index-period equals
$\sum_{\rho=1}^{\delta-1}c(\rho).$
@@ 5. Some numerical examples
This section illustrates the arithmetic mechanism of Section 4 with three examples. The first shows the degenerate situation $d=1$, where every $(a+1)$-gap yields exactly one collision window. The second gives an explicit computation in the first non-trivial case $d=2$. The third demonstrates how the parameters $L$ and $R$ determine $c(\rho)$.
@@
The gaps of size $a+1=4$ occur exactly $\delta-1=3$ times. Since $d=1$, a collision window consists solely of a single $(a+1)$-gap and no additional $1$-gaps. Thus each such gap contributes exactly one collision window, so $c(\rho)=1$ for all $\rho$ and
$\sum_{\rho}c(\rho)=3=ad.$
@@
Since $d=2$, each collision window must contain one $(a+1)$-gap and one neighbouring $1$-gap. Inspecting the sequence shows that some six-gaps are adjacent to two $1$-gaps while others are adjacent to only one. Writing $c(\rho)$ for the number of collision windows containing the $(a+1)$-gap at index $\rho$, we obtain
@@
Hence among the $\delta-1=6$ many $(a+1)$-gaps, exactly two have $c(\rho)=1$ and the remaining four have $c(\rho)=2$. Therefore
$\sum_{\rho}c(\rho)=1+1+2+2+2+2=10=ad.$
@@
Let $a=4$ and $\delta=7$. Then $d=3$, so each collision window must contain exactly one gap of size $a+1=5$ and exactly d1=2d-1=2 gaps of size $1$. This example illustrates how the parameters $L$ and $R$ determine the number $c(\rho)$. The index-period has length $a\delta=28$. The gap formula reads
@@
Consider the gap corresponding to $n+1=12$. Writing $12=qa$ with $q=3$ and reducing modulo $\delta=7$ gives $12\equiv 5\pmod{7}$, so $\rho=5$. We now compute
$L=\min(a-1,\rho-1)=\min(3,4)=3,$
@@
force $x=1$ or $x=2$. Hence this $(a+1)$-gap contributes $c=2$ collision windows. This agrees with the right-restricted case of the general formula (where $a\leqslant\rho\leqslant\delta-1$).
6. The additive subtraction nimbers finally find their way home
@@
In one heap-period, $|\mathcal{W}_{3}|_{h}=a^{2}$.
@@
By Theorem 9, we have that $|C|_{h}=ad$, where $d=\delta-a$ and consequently
$|\mathcal{W}_{3}|_{h}$ $=|W_{0}-\delta|_{h}-|C|_{h}$
@@
$\mathbb{N}=\mathcal{W}_{0}\sqcup\mathcal{W}_{1}\sqcup\mathcal{W}_{2}\sqcup\mathcal{W}_{3},$
@@
We have observed $|\mathcal{W}_{0}|_{h}=|\mathcal{W}_{1}|_{h}=|\mathcal{W}_{2}|_{h}=a\delta$, and, via Corollary 17, $|\mathcal{W}_{3}|_{h}=a^{2}$. In total, the disjoint union
@@
$|\mathcal{W}_{0}|_{h}+|\mathcal{W}_{1}|_{h}+|\mathcal{W}_{2}|_{h}+|\mathcal{W}_{3}|_{h}=3a\delta+a^{2}$
positions, exactly matching the heap-period length in the primitive quadratic regime. Recall, the index-period is $a\delta$, and since $a\delta$ is simultaneously a multiple of $a$ and $\delta$, the direct computation in (3) gives the corresponding period length $w_{n+a\delta}-w_{n}=(3\delta+a)a$.